# Quasi-Deterministic Modeling of Sub-THz Band Access Channels in Street Canyon Environments

### v1

@@ VI-A Path Loss
(15) where, is the Tx-Rx separation distance, is the path loss exponent (PLE), is the center frequency, and is the speed of light in .
@@
The results of fitting the omnidirectional PLs for both LoS and NLoS scenarios using the aforementioned models are shown in Fig.13, and the corresponding model parameters are summarized in TableIV as well. From the fitting results, it can be observed that in the LoS scenario, both the CI and FI models yield similar parameter values, such as the PLE ( or ) and the standard deviation (), for each frequency band. In contrast, under the NLoS scenario, the values of the FI model for both the GHz and GHz bands are smaller than those of the CI model, suggesting that the FI model provides a better fit for NLoS conditions in this environment. Nonetheless, it is premature to draw definitive conclusions due to the limited number of samples.
From Fig. 13(a), it can be observed that the measured PL of each frequency band has distinctly exhibited distance dependency and several dB improvement can be achieved over FSPL due to multipath power in LoS scenario. Furthermore, when comparing the measured PL with the FSPL in each band, the difference at GHz is slightly smaller than that at GHz. This indicates that the GHz band undergoes somewhat greater degradation, which is consistent with the cluster characteristics discussed in the previous section.
In the NLoS scenario, the PLs are significantly higher, with a large shadowing distribution. This is evident from the standard deviation () values of the fitting parameters for both models, as shown in Fig. 13(b). At all positions, the PLs are considerably larger than the FSPL.
VI-B Large Scale Channel Parameters (LSPs)
The PDP and PAP are calculated using (1) and (3), respectively. The obtained power profiles are then processed according to an acceptance criterion. In particular, only those samples are retained that are either within dB of the peak sample in the power profiles [22] or within the cutoff level, whichever is higher. Similarly, the corresponding PAP is processed according to the same acceptance criterion before being used to evaluate the respective angular spreads (AS), namely the Azimuth-of-Departure (ASD) and the Azimuth-of-Arrival (ASA).
The temporal dispersion, on the other hand, is evaluated as the second moment of the PDP as follows
@@ VI-B Large Scale Channel Parameters (LSPs)
Similarly, ASD at Tx and ASA at Rx are obtained in terms of the relevant PAP as
@@
The power content in the LoS cluster compared to other significant MPCs is quantified as Rician KF and can be evaluated as
@@
The CDF plots of DS, ASD, ASA, and Rician KF for the considered scenarios are presented in Fig. 14. Based on the analysis, the key observations can be summarized as follows:
Due to several outlier MPCs with large delays observed at specific Rx positions, the DS exhibits large deviations above the 80% level; nevertheless, it remains similarly distributed at both frequencies up to this percentile.
The average ASA and ASD are approximately dB and dB, respectively, in the LoS scenario. They are similarly distributed at both frequencies; however, the ASA is consistently larger than the ASD, which can be attributed to the proximity of the Rx to the SW.
The K-factors range from to dB, with an average of to dB. No significant difference is observed between the distributions at GHz and GHz.
@@ VII Conclusion
This paper presented an extensive D-D channel measurement campaign in a street-canyon environment at GHz and GHz, under both LoS and NLoS conditions, using an in-house-developed dual-band channel sounder. By applying clustering to merged datasets across the two frequency bands, both common and distinct multipath clusters were identified, enabling a comparative analysis of their frequency-dependent characteristics. The results show that the dominant scattering processes are primarily the direct path and single-bounce reflections from the north and south walls. Statistical analyses reveal no significant differences, but slightly greater degradation is observed at GHz.
A QD channel model was proposed, integrating deterministic components, LoS and major reflections from the north and south walls, with random components representing higher-order or irregular multipath contributions. Furthermore, LSPs, including PL, DS, AS, and the Rician KF, were characterized. These findings deepen the understanding of sub-THz propagation in urban street canyons and provide essential inputs for the development of accurate, frequency-dependent channel models, supporting the design and deployment of future 6G systems.

### v2

@@ V-A Path Loss
(10) where, is the Tx-Rx separation distance, is the path loss exponent (PLE), is the center frequency, and is the speed of light in . denotes the log-normal shadow-fading term with zero mean and standard deviation .
@@
| Scen. | Freq. | CI Model | FI Model |
| LoS | GHz | ||
| GHz | |||
| NLoS∗ | GHz | ||
| GHz | |||
| ∗Preliminary values due to limited sample size. | |||
The results of fitting the omnidirectional PLs for both LoS and NLoS scenarios using the aforementioned models are shown in Fig. 9, and the corresponding model parameters are summarized in Table II. For comparison, the latest ITU-R P.1411-13 site-general model (ABG model), which supports frequencies up to 300 GHz, is also plotted as a reference.
ABG model:
(12) where , , and are the distance slope, intercept, and frequency slope, respectively.
From the fitting results, it can be observed that in the LoS scenario, both the CI and FI models yield similar parameter values, such as the PLE ( or ) and the standard deviation (), for each frequency band. In contrast, under the NLoS scenario, the values of the FI model for both the GHz and GHz bands are smaller than those of the CI model, suggesting that the FI model provides a better fit for NLoS conditions in this environment. Nonetheless, it is premature to draw definitive conclusions due to the limited number of samples.
From Fig. 9(a), it can be observed that the measured PL of each frequency band has distinctly exhibited distance dependency and several dB improvement can be achieved over FSPL due to multipath power in the LoS scenario. Furthermore, when comparing the measured PL with the FSPL in each band, the difference at GHz is slightly smaller than that at GHz. This indicates that the GHz band undergoes somewhat greater degradation, which is consistent with the cluster characteristics discussed in the previous section. On the other hand, the ITU-R model tends to overestimate the PL compared to the measured data, yielding values that exceed the FSPL. This discrepancy is likely because the ITU-R model is fitted using a vast dataset primarily from lower frequency bands, which may limit its prediction accuracy at sub-THz frequencies in this specific environment.
In the NLoS scenario, the PLs are significantly higher than the FSPL, exhibiting a large shadowing distribution as evidenced by of the fitting parameters shown in Fig. 9(b). However, due to the limited sample size in this scenario, these parameters should be considered preliminary and indicative rather than definitive. The measured PL exhibits a milder slope than the ITU-R predictions. This deviation highlights the site-specific nature of the environmental geometry.
Relating our results to prior outdoor sub-THz measurements using the same CI model with a 1 m reference distance, our omnidirectional PLEs at GHz closely match the GHz UMi results reported in [7] (LoS , NLoS ). Overall, our near-free-space LoS slopes are consistent with street/campus measurements in the – GHz range [6, 8, 9], while remaining differences (especially in NLoS) are mainly driven by site geometry and measurement/processing method.
V-B Large Scale Channel Parameters (LSPs)
The PDP and PAP are calculated using (1) and (3), respectively. The obtained power profiles are then processed according to an acceptance criterion. In particular, only those samples are retained that are either within dB of the peak sample in the power profiles [20] or within the cutoff level, whichever is higher. The temporal dispersion, RMS DS, is evaluated as the second moment of the PDP as follows
@@ V-B Large Scale Channel Parameters (LSPs)
Similarly, the corresponding PAP is processed according to the same acceptance criterion before being used to evaluate the respective angular spreads (AS). Specifically, the AS of Departure (ASD) at the Tx and the AS of Arrival (ASA) at the Rx are calculated based on the definition in 3GPP TR 38.901 v18.0.0 [2, Annex A], expressed as
@@
| Param. | Model | LoS | NLoS∗ | |||
| GHz | GHz | GHz | GHz | |||
| DS | Meas. | |||||
| 3GPP | ||||||
| ASD | Meas. | |||||
| 3GPP | ||||||
| ASA | Meas. | |||||
| 3GPP | ||||||
| -factor [dB] | Meas. | — | — | |||
| 3GPP | — | — | ||||
| ∗Preliminary values due to limited sample size. | ||||||
The power content in the LoS cluster compared to other significant MPCs is quantified as Rician -factor and can be evaluated as
@@
The CDF plots of DS, ASD, ASA, and Rician -factor for the considered scenarios are presented in Fig. 10. Based on the analysis, the key observations can be summarized as follows:
The average DS is approximately ns at GHz and ns at GHz. Compared to the 3GPP extrapolation, the measured DS is generally smaller, indicating that the sub-THz channel in this environment is less dispersive in the delay domain than the standard model predicts.
In the LoS scenario, the average ASD is approximately at GHz and at GHz, while the average ASA is and , respectively. While the distributions remain statistically similar across the two frequencies, a slight reduction is observed at GHz. The ASD is consistently larger than the ASA, which can be attributed to the proximity of the Rx to the street walls (SW). Compared to the 3GPP model, the measured ASD shows reasonable agreement, whereas the measured ASA is notably smaller than the prediction. This discrepancy highlights the site-specific nature of the environment as well as the multipath sparsity characteristic of sub-THz bands.
The -factors range from to dB, with an average of dB at GHz and dB at GHz (a slight increase at GHz). No significant difference is observed between the distributions at the two frequencies. These results align well with the 3GPP model, which predicts a mean -factor of dB, confirming that the dominance of the direct path remains consistent with lower-frequency extrapolations.
To statistically verify the frequency dependence, we conducted t-tests [23] on the extracted log-valued LSPs. The results indicate no statistically significant differences (-value, ) between GHz and GHz for any parameter. Specifically, the Rician -factor demonstrated significant similarity (), while the dispersion parameters, DS (), ASD (), and ASA (), also showed no significant shifts.However, their mean values did exhibit consistent physical trends: a slight increase in the -factor and a slight decrease in dispersion parameters at GHz. These minor variations are attributed to increased interaction losses (e.g., rough surface scattering), which tend to suppress higher-order MPCs more severely than the LoS path. Consequently, this statistical evidence confirms that the channel’s temporal and spatial dispersion is primarily governed by the site geometry rather than frequency-dependent scattering mechanisms.
The measurement configuration (BS height: m, and UE height: m) represents a low-height access link in a street-canyon environment (lamppost-installed BS to UE). The reported PL and LSPs are derived from the measurement distance ranges only (LoS: – m, and NLoS: up to m). While the LoS dominance and multipath sparsity are expected to hold for similar street-canyon scenarios, the channel parameters may vary with street geometry/materials and, in particular, with Tx height due to changes in elevation geometry and the visibility/strength of wall and ground interactions.
VI Channel modeling
Based on the investigation of the propagation mechanisms, this section proposes a Q-D channel model inspired by those developed for corridor scenarios [13]. The discussion begins with the mathematical representation of the channel, followed by the modeling and generation of the deterministic and random components that constitute the channel. It should be noted that the channel model is developed only for the LoS scenario, due to the limited amount of available data.
VI-A Q-D channel model
Due to the short wavelength of sub-THz waves, specular reflection dominates and can be effectively predicted using deterministic methods such as ray tracing. Building on this characteristic as the deterministic component, a Q-D channel model [5, 12] further accounts for non-negligible interactions arising from higher-order reflections and scattering caused by object shape and surface roughness, modeling them as random components. Namely, the propagation channel can be modeled as
where the deterministic components, , and random components, are separately represented. Note that diffuse scattering is modeled as an additional contribution associated with both the deterministic and random components (i.e., as a local micro-cluster around each component).
where ‘’ and ‘’ denote the single-bounce reflections by NW and SW, respectively. denotes a random variable that captures the randomness of the SB path existence and, in principle, its gain variations (e.g., due to surface roughness and associated diffuse scattering). In this work, however, is modeled using a two-state Markov process to represent the presence/absence of the SB path only, while a comprehensive stochastic modeling of the SB gain fluctuations (i.e., based on Fig.11(b)) is left for future work.
| Freq. | Transition probability matrix |
| GHz | , |
| GHz | , |
| Freq. | 1/ | |||
| GHz | ||||
| GHz |
VI-B Deterministic components
Fig. 11(a) illustrates the distribution of the differences between the measured and calculated powers for the SB reflections from the NW and SW at both frequencies, where missing Rx positions indicate the absence of a signal. Fig.11(b) further presents the empirical CDFs together with their fitted models. To capture this probabilistic behavior induced by factors such as surface roughness, these models can be combined with a two-state Markov chain to represent signal presence or absence, as proposed in [13, 24]. This property of the Markov chain is represented by the transition probability matrix where is the probability of transitioning from ‘Absent’ to ‘Present’, and is the probability from ‘Present’ to ‘Absent’. The modeling of path appearance is carried out separately for the SB reflections from the NW and SW at each frequency. Table IV summarizes the estimated transition probability matrices for each reflector. The steady-state probability of path absence, derived from the element of the limiting matrix , increase from GHz to GHz. Specifically, the absence probability () rises from to for the NW reflection, and from to for the SW reflection. This trend is driven by greater scattering loss and diffuseness at GHz.
VI-C Random components
As described above, non-negligible random interactions should be incorporated into the channel model for a more accurate representation. Fig. 12 shows the distributions of relative PG for the deterministic and random components, where the SB reflections from the NW and SW are treated as deterministic components and all other components are treated as random. Note that this partition is adopted for the LoS scenario. For NLoS conditions, a dominant DB mechanism (e.g., SWNW in Fig. 6(b)) may be incorporated as an additional deterministic component. Evidently, the random components make a significant contribution to the overall power. As explicated earlier, MPCs originating from non-predominant higher-order reflections and difficult-to-identify propagation paths are modeled as random components.
and has been fitted accordingly in Fig. 13(a). Here, represents the arrival time of the th MPCs, and is the reciprocal of the mean interarrival time (arrival rate). The fitting results presented in Table V show that the mean interarrival time at GHz is ns, which is higher than ns at GHz. It should be noted that these values are more than an order of magnitude larger than those observed in indoor corridor environments, owing to the outdoor setting.
where , represents the slope, and represents the intercept. Furthermore, and represent the delay and path gain of the LoS component at distance , respectively. denotes a zero-mean Gaussian random variable with variance , representing shadow fading. The parameters and are estimated by fitting the plot in Fig.13(b) using the maximum likelihood method [25], considering a noise level of dB. The results yield and at GHz, and and at GHz, illustrating a similar trend at both frequencies. Furthermore, as shown in Fig. 13(c), the path gain deviation to the model expressed in dB fits a normal distribution with a standard deviations of at GHz and at GHz, respectively. The obtained parameters for random cluster model are tabulated in Table V.
VI-D Quantitative Validation of the Proposed Model
To assess the accuracy of the proposed model prediction, we quantitatively validate the generated channels against the measurement data.
VI-D1 PL accuracy
We first evaluate the PL prediction accuracy by comparing the measured and synthesized PL over all receiver locations. For each receiver position, the model generates multiple channel realizations to account for the stochastic components. Fig. 15 shows that the synthesized PL closely follows the measured distance-dependent trend at both GHz and GHz. The RMS error (RMSE) between measured and modeled PL remains small (approximately dB or less), confirming that the proposed model reliably captures the dominant behavior governed by site geometry.
VI-D2 Statistical consistency of LSPs
Next, we examine whether the model reproduces the statistical distributions of representative LSPs. Fig. 16 illustrates the validation results for DS and -factor, displaying both the spatial variability (scatter plots) and the CDFs. We applied the two-sample Kolmogorov–Smirnov (KS) test to quantify the goodness of fit [26]. As indicated, the KS statistics are consistently below the critical value, and the -values exceed the significance level of in all cases. Consequently, the null hypothesis that the modeled and measured LSPs are drawn from the same distribution cannot be rejected.
The close overlap between the measured and modeled CDFs, combined with the favorable KS-test results, demonstrates that the proposed Q-D model accurately reproduces not only the average trends but also the variability of the dispersion characteristics. Importantly, the model effectively captures the frequency-dependent sparsity observed at GHz, reflecting the reduced multipath richness and weaker dispersion compared to the GHz band.
@@ VII Conclusion
This paper presented an extensive D-D channel measurement campaign in a street-canyon environment at GHz and GHz, under both LoS and NLoS conditions, using an in-house-developed dual-band channel sounder. By applying joint clustering to merged datasets, we identified common multipath clusters and performed a comparative analysis of their frequency-dependent behaviors. The results show that the dominant propagation mechanisms are primarily the direct path and single-bounce reflections from the north and south walls. While statistical analyses of LSPs revealed no significant differences between the bands indicating geometric dominance, slightly greater interaction loss was observed at GHz. This is attributed to increased interaction losses (e.g., rough surface scattering), which tend to suppress higher-order MPCs more severely than the LoS path.
A Q-D channel model was proposed, integrating deterministic components, specifically the LoS and major reflections from the north and south walls, with random components representing higher-order or irregular multipath contributions. Furthermore, LSPs, including PL, DS, AS, and the Rician -factor, were characterized. These contributions deepen the understanding of sub-THz propagation in urban street canyons and provide essential inputs for developing accurate, frequency-dependent channel models for future 6G systems.